Circular No. 8821

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2467 CYGNI = NOVA CYGNI 2007 S. Nakano, Sumoto, Japan, reports the discovery by Akihiko Tago

(Tsuyama, Okayama-ken) of a possible nova (mag 7.4) on three 30-s CCD exposures taken around Mar. 15.787 UT with a Canon 20Da digital camera (+ 70-mm f/3.2 lens; limiting magnitude 13.2 ± 0.2), with Tago giving the position of the variable as α = 20^{h}28^{m}15^{s}, δ = +41^{o}49′.0 (equinox 2000.0), and noting that nothing was visible at this position on his frame taken on Mar. 12.796 (limiting mag 12). Tago adds that an image taken by himself on Mar. 16.75 shows the variable at mag 7.5. Nakano provides his own measure of the position end figures (estimated uncertainty perhaps ± 5′′ or more) from Tago’s discovery image: 13^{s}.22, 48′45′′.5, magnitude 6.8 ± 0.4. Nakano further adds that K. Nishiyama and T. Sakamoto, Bisei Spaceguard Center, have confirmed the variable star at mag 6.7 on an unfiltered CCD image taken with the 1.0-m f/3.0 reflector on Mar. 16.768, giving the object’s position end figures as 12^{s}.52, 48′36′′.5. A red USNO-catalogued star is located at position end figures 12^{s}.505, 48′36′′.69 (red mag 18.5).

K. Ayani, Bisei Astronomical Observatory (BAO), writes that lowresolution spectra (range 380-800 nm; resolution 0.55 nm at Hα), obtained with the BAO 1.01-m telescope on Mar. 16.8 UT, show that this variable is a nova. The spectra show Balmer and Fe II absorption lines on a red continuum. The Hα line has a P-Cyg profile whose absorption bottom is blueshifted by 1200 km/s with respect to the emission peak.

H. Naito and M. Sakamoto, Nishi-Harima Astronomical Observatory, report that they obtained a low-resolution spectrogram (range 410-670 nm; resolution 2000 at Hα) of the nova on Mar. 16.792 UT with the 2.0-m NAYUTA telescope (+ MALLS). The spectrum shows a broad Hα line with a P-Cyg profile and Fe II absorption lines, which suggests that the variable is indeed a classical nova near maximum light. A narrow Na I D absorption feature suggests the existence of a considerable interstellar reddening. The FWHM of the Hα emission is 950 km/s, and its absorption minimum is blueshifted by 1150 km/s from the emission peak (measured by a Gaussian fitting).

Visual magnitude estimates: Mar. 17.010 UT, 7.9 (K. Hornoch, Lelekovice, Czech Republic); 17.146, 8.1 (K. Sárneczky, Budapest, Hungary); 18.453, 7.7 (J. D. West, Mulvane, KS, U.S.A.); 19.012, 8.5 (P. Schmeer, Bischmisheim, Germany).

N. N. Samus, Institute of Astronomy, Russian Academy of Sciences, informs us that this nova has been given the designation V2467 Cyg.

2007 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green